40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Entry  Tue Oct 9 00:44:22 2018, johannes, Notes, Cryo ISS, New Transmission PDs and ISS 8x
    Reply  Thu Oct 11 01:25:43 2018, johannes, Notes, Cryo ISS, ISS Servo Characterization 7x
       Reply  Tue Oct 16 00:02:15 2018, johannes, Notes, Cryo ISS, ISS Servo Characterization planex_rin.pdf
Message ID: 2180     Entry time: Tue Oct 9 00:44:22 2018     Reply to this: 2181
Author: johannes 
Type: Notes 
Category: Cryo ISS 
Subject: New Transmission PDs and ISS 

I assembled one of Anchal's revision of Zach's M2ISS transimpedance amplifier for shot noise limited measurements of the transmitted light power. To keep it simple I did not populate the switching logic for switchable transimpedance gain, and also omitted the whitening stage. The 5V bias voltage for the PDs is buffered by individual OP27. Since the photocurrent will be no larger than ~4mA I omitted the BUF634 on the bias side. I did add BUF634 in the transimpedance stage so the circuit would be able to drive 50Ohm loads if needed.

I was using an LT1128 at first, as it should deliver the lowest nominal noise floor for a transimpedance gain of 1kOhm, but I was observing oscillations at ~1.2 MHz which I couldn't suppress without seriously affecting the bandwidth of the detector. Koji pointed out that the LT1128's open loop gain has small phase margin at this frequency, as visible in the data sheet

I replaced the LT1128 with an OPA827 which performs only marginally worse according to Disha's LISO simulations. This remedied the oscillations. I still had to put 22pF in parallel with the 1kOhm transimpedance resistor for loop stability. I boxed the satellite amplifier and mounted two 3mm InGaAs diodes in 4" mounts that I had extruded from Zach's M2ISS design. Using an SR560 I measured the PD noise, which is shown in Attachment 5. I need to check why the west path is worse at low and high frequencies, but the baseline NEP for both is about 6 pW/rtHz, pretty much on point with Disha's LISO simulation.

I changed the optical configuration of the experiment back to sending the undiffracted beams from the AOMs into the cavities. Attachment 6 shows the transfer function from input to the commercial AOM driver to output of the new custom transmission PDs.

I also completely modded the existing ISS box to host two of Anchal's new ISS servo. I finished debugging the circuits and now both are doing exactly as asked - noise measurement pending. Attachment 7 shows the open loop gain of just the servo boards, and Attachment 8 the product of servo TF and plant TF. I haven't yet closed the loop, but it looks like some external attenuation would be enough to make the system stable. I'll try this tomorrow.

 

 

Attachment 1: 20181005_164259.jpg  3.306 MB  Uploaded Tue Oct 9 01:45:08 2018  | Hide | Hide all
20181005_164259.jpg
Attachment 2: 20181005_164652.jpg  2.735 MB  Uploaded Tue Oct 9 01:45:23 2018  | Hide | Hide all
20181005_164652.jpg
Attachment 3: 20180821_190651.jpg  3.743 MB  Uploaded Tue Oct 9 01:48:14 2018  | Hide | Hide all
20180821_190651.jpg
Attachment 5: TPD_noise.pdf  78 kB  Uploaded Tue Oct 9 02:05:09 2018  | Hide | Hide all
TPD_noise.pdf
Attachment 6: iss_plant.pdf  25 kB  Uploaded Tue Oct 9 02:08:48 2018  | Hide | Hide all
iss_plant.pdf
Attachment 7: ISS_Servo.pdf  23 kB  Uploaded Tue Oct 9 02:14:13 2018  | Hide | Hide all
ISS_Servo.pdf
Attachment 8: ISS_OLG.pdf  28 kB  Uploaded Tue Oct 9 02:15:13 2018  | Hide | Hide all
ISS_OLG.pdf
ELOG V3.1.3-